OpenStudy (anonymous):

I have a step in a question where tanx=-1 then tan^-1(-1)=-pi/4. how did that get to tan (pi - pi/4), tan(2pi - pi/4)?

OpenStudy (jhannybean):

If you're trying to find the general solutions of x in a radian circle for each interval of \(\dfrac{\pi}{4}\), then you have to consider \(\theta\) goes from \(0 \rightarrow 2\pi\)

OpenStudy (anonymous):

I was trying to solve for x over the interval 0<=x<=2pi. I'm not too sure how to use the radian circle for this

OpenStudy (jhannybean):

Look for all the possible radians that give you -1 as a solution around the unit circle.